OpenStudy (anonymous):

Are these lines parallel or perpendicular y=5x+3 and y=-5x+8

OpenStudy (anonymous):

When they are perpendicular, its only if when you multiply their gradients, they must give you -1. in this case 5 multiply -5=-25. therefore they are not perpedicular

OpenStudy (anonymous):

So they they are parallel

OpenStudy (anonymous):

not even parallel.

OpenStudy (anonymous):

Two lines are parallel if they have the same gradent. Example. The lines y = 2x + 1 and y = 2x + 3 are parallel, because both have a gradient of 2.

OpenStudy (anonymous):

What about these lines x/3-4 and 1/3x+2

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

Parallel???

OpenStudy (anonymous):

Yes , but why?

OpenStudy (anonymous):

Because they have the same gradent

OpenStudy (anonymous):

\[y=\frac{ x }{ 3 }-4\] and \[y=\frac{ 1 }{ 3 }x +2\] ?

OpenStudy (anonymous):

Dats correct if dats the equations i typed above

OpenStudy (anonymous):

wanderful

OpenStudy (anonymous):

Alright thank you. And how do I find the x intercept on this problem 3x+4y=8

OpenStudy (anonymous):

an x-intercept is a point on the graph where y is zero, and a y-intercept is a point on the graph where x is zero.

OpenStudy (anonymous):

But how do I figure thst out wit that equation

OpenStudy (anonymous):

Using the definitions of the intercepts, you should proceed as follows: x-intercept(s): y = 0 for the x-intercept(s), so: 3x+4(0)= 8 and solve for x

OpenStudy (anonymous):

But how??

OpenStudy (anonymous):

Oh sorry... \[3x+4y=8\] \[ x-intercept, y=0\] where you see y, you subtitutes 0 therefore we have \[3x+4(0)=8\] \[3x+0=8\] \[3x=8\] \[x=\frac{ 8 }{ 3 }\] \[\therefore (\frac{ 8 }{ 3 };0)\]